If you have 4 vegetables, 3 meats, and 4 fruits. How many combinations can you get?
Answer by jim_thompson5910(35256)

You will have 4*3*4 = 12*4 = 48 different combinations (assuming you can only pick 1 veggie, 1 meat and 1 fruit)